# Uncertainty-Aware Probabilistic Graph Neural Networks for Road-Level Traffic Accident Prediction

### v3

@@ 4.6.1 Temporal Stability of Metrics
Evaluating temporal stability is crucial for evaluating a model’s performance over multiple prediction time steps, particularly in data-driven methods, where superior results in specific instances might mask potential underfitting issues (Ding et al.,, 2022; Mannering,, 2018). Our primary goal is to achieve consistently low uncertainty, as indicated by the MPIW metric in three boroughs, depicted in Figure 5. We strive to maintain a narrow yet comprehensive MPIW throughout the prediction steps to ensure reliable prediction intervals for traffic accidents.
@@
As indicated in Figure 5, the MPIW is typically lowest on the first day of testing and generally increases over time. However, this increase is smoother in our STZITD-GNNs model compared to the more erratic fluctuations observed in baseline models. This smoother increase reflects our model’s ability to handle the inherent uncertainties and the variability of risk levels, which might be less accurately captured by models with misspecified distributions. Moreover, boroughs like Westminster and Lambeth, known for higher instances of zero-accident risk values, show more fluctuating MPIWs, suggesting greater uncertainties in prediction intervals due to the sparsity of traffic accidents. Despite these challenges, our model consistently delivers stable and accurate predictions across different boroughs.
In the evaluation of the temporal stability for the estimation of points, depicted in Figure 6, the STZITD-GNN customised approach emphasises the advantages of applying the most appropriate probabilistic assumptions, enhancing spatiotemporal representations, and managing the various aspects of traffic accidents. It is apparent that the STG-GNN model consistently shows the weakest performance across all temporal prediction periods, with deterministic models also performing poorly. Furthermore, models based on the Negative Binomial (NB) distribution exhibit unsatisfactory performance characterised by higher fluctuation and higher error rates. This inadequacy stems from the focus of NB models on binary accident-occurrence outcomes, which often overlook the complexities of long-tail accident risk values. In contrast, our model exhibits significantly less variance across all three evaluated metrics at each time step, demonstrating stability even toward the end of the prediction horizon. This robust performance is attributed to the model’s ability to dynamically fine-tune prediction intervals using four critical parameters, effectively addressing the challenges posed by the sparsity and skewed distribution of traffic accident data.
@@
The final component of the evaluation focusses on quantifying both non-accident roads and high-risk roads where real accidents occurred, as shown in Figure 7. It is evident that the STZITD-GNN model consistently excels in both metrics, maintaining stable performance with less than 18% variance throughout the 14-step period. As illustrated in Figure 7(a), deterministic models exhibit less variability but face substantial challenges in accurately identifying true nonaccident roads, often misclassifying all roads as having no accident risk in extreme zero inflation scenarios, leading to overfitting. Conversely, probabilistic models employing Gaussian distributions struggle to accommodate the high prevalence of zero values, resulting in poor performance and high sensitivity in the ZR metrics over time. This issue arises because each time step may present unique scenarios of zero-valued roads, which Gaussian models often misinterpret due to their uniform mean and variance assumptions.
In the identification of High-Risk Accident Roads as illustrated in Figure 7(b), the ZITD distribution particularly excels. It incorporates a weighted dispersion parameter ($\phi$) to distinguish the lower accident risk from the non-accident road, displaying the best of low, middle and high risk roads with long-tailed distributions. This adaptability to the complex temporal dynamics of road-level accident data at each time step not only mitigates overfitting, but also stabilises the model performance, enabling it to accurately capture real accident occurrences in critical geographic locations.
@@
(b) Accuate Hit at Top 20% Risky Roads over Time Performance
Figure 7: Time-wise Comparison for the Non Accident Roads Quantification and Accurate Accident Hit Metrics
4.6.2 Analysis of Long-tail Accident Data Characteristics
To explore the impact of inherently long-tailed data on our model, we created a three-dimensional surface plot over four information sources, as illustrated in Figure 8. Due to visualisation constraints, it is not feasible to represent all four parameters in a single plot. We therefore focus on the TD parameters $\phi$, $\rho$, and $\mu$, with the understanding that the sparsity parameter $\pi$ significantly influences the original TD parameters. This plot illustrates the predicted accident values, with the actual accident values $y$ represented by the colour gradient.
A distinct trend is evident in the learnt values of $\rho$, consistently approaching 2, suggesting a preference for more skewed distributions as $\rho$ increases from 1 to 2 (Tweedie et al.,, 1984; Kurz,, 2017). This trend corresponds with the long-tailed nature of our data and validates the probabilistic assumptions. The increase in $\rho$ serves as an adaptive response to the component of the loss function $y\frac{\mu^{1-\rho}}{1-\rho}(y>0)$, effectively penalising the underestimation of true, although infrequent, higher values of accident risk within the long tail. This adjustment directs the model to focus more on such kinds of traffic accident rates, continually benefiting from significant enhancements in model sensitivity and accuracy.
Additionally, significant variations in the values $\phi$ are observed, particularly concerning data with zero value, which serve as the dispersion parameter. For example, in Westminster, $\phi$ is significantly lower than in other boroughs, indicating fewer zero-inflation issues. Traditionally, $\phi$ in the TD distribution is crucial for managing the frequency of zero occurrences to effectively address zero inflation. However, due to the highly imbalanced nature of our data, a single $\phi$ parameter often fails to capture the prevalence of zeros, as demonstrated by the overlap of plots in areas of lower and zero accident values. In our ZITD model, an enhanced $\phi$, augmented by the sparsity parameter $\pi$, manages these features, offering a precise depiction of the inherent sparsity and discrete uncertainty in road-level accident predictions.
Moreover, the model parameter $\mu$, which denotes the mean of the predicted accident values, shows a strong correlation with the actual accident values $y$, as evidenced by the colour alignment in the graphs. This correlation not only substantiates the effectiveness of the model in estimating mean points, but also affirms its capability to accurately mirror and adapt to the complexities of urban accident data.
@@ 4.6.2 Analysis of Long-tail Accident Data Characteristics
Figure 8: Three Tweedie parameter visualization on Lambeth, Tower Hamlet, and Westminster with multi-step predict. The colour bar shows the predicted accident from the highest to zero.
@@ 5 Conclusion and Discussion
This paper introduces STZITD-GNN, a probabilistic deep learning framework that, to the best of our knowledge, is the first uncertainty-aware model specifically tailored for road-level accident prediction. Developed from data-driven insights, the STZITD-GNN utilises predefined distribution characteristics of traffic accident data to enhance the representation of sparse traffic accident scenarios. It effectively detects finer-level correlations in traffic accidents, thereby delivering precise predictions with narrow prediction intervals and reduced uncertainties. Through systematic multi-timescale experiments conducted on three real-life urban accident datasets, our results demonstrate that the efficacy of our STZITD-GNN approach surpasses that of state-of-the-art probabilistic deep learning models as well as deterministic models. The key contributions of this study are summarised as follows.
@@
  • Innovative Distribution Extension: We have adapted the conventional TD to a ZITD distribution, introducing a sparsity parameter, $\pi$, integrated with the dispersion parameter $\phi$. This modification tailors the model to our data-driven assumptions, representing the dispersed and extreme zero-inflated nature of accident data at the road level.
@@
  • Integration with ST-GNN: The incorporation of the ZITD encoder into the ST-GNN framework makes the STZITD-GNN model an end-to-end framework to handle the complex spatiotemporal dynamics of accidents. It surpasses traditional point estimation methods by enabling the mapping of output variance and prediction intervals. Additionally, the model facilitates an interpretable, uncertainty-aware process that elucidates the underlying factors contributing to improved prediction accuracy.
@@
  • Empirical Validation Across Boroughs:The model has shown improved predictive accuracy and minimal uncertainties compared with benchmark models in three London boroughs, each featuring unique spatial accident patterns in multi-step scenarios. This performance confirms the robustness and generalizability of the model across diverse urban settings, even with sparse datasets. Additionally, the model’s stable performance across heterogeneous datasets, which vary in accident determination criteria, further underscores its reliability and effectiveness in urban traffic risk assessment.
@@
It should be recognised that the challenges of overdispersion and zero inflation extend beyond traffic accidents, affecting a wide array of urban issues such as crime rates. This universality suggests that our proposed model could be applicable to a broader range of contexts beyond its current usage.
@@
While our research demonstrates its effectiveness in predicting traffic accidents, there are several limitations that provide direction for future research. We aim to enhance the accuracy of mapping accidents at finer temporal resolutions and plan to expand the model’s applicability to various countries under consistent assumptions. Future developments will focus on creating a more adaptable model that can accommodate geographic differences, thus increasing its usability across diverse regions. In addition, we intend to integrate more diverse data types into our model. Incorporating GPS data, for example, could provide more detailed insight into traffic exposure. To further improve the accuracy and utility of our model, we also plan to use more comprehensive data sets, including urban characteristics derived from street or satellite images. This would enrich our model’s contextual understanding and could identify factors that exacerbate traffic accident impacts across different locations and times. Addressing these challenges in future studies will improve the robustness and applicability of the framework in real-world scenarios, ultimately contributing to the development of more effective traffic management systems and accident prevention strategies.
@@
In conclusion, the development of the STZITD-GNN model marks a significant advancement in the prediction of urban road-level traffic accidents. However, this is just the beginning of an ongoing effort to refine and enhance this model. By integrating additional data sources and continually improving our methodologies, our goal is to develop a tool that not only mitigates traffic risks but also significantly contributes to global urban safety strategies.
@@ Appendix A Tweedie Table
Appendix B Temporal encoder – GRU
Traffic accidents demonstrate significant temporal correlation, underscoring the idea that the dynamic temporal features of roads encapsulate a wealth of information. This observation has been acknowledged and exploited in prior methodologies (Wang et al., 2021a, ; Zhao et al.,, 2019). As shown in Figure 1, to seize these temporal proximity features, we implement a GRU (Chung et al.,, 2014). The GRU, in comparison to traditional LSTM networks, offers the advantage of learning relatively long-term dependencies without facing issues of vanishing and exploding gradients. In the GRU framework, there isn’t a segregation of internal and external states, as seen in LSTM networks. Instead, it addresses the gradient vanishing and exploding problem by directly integrating a linear dependency between the current network state $h_{t}$ and the preceding state $h_{t-1}$. Consequently, while a GRU preserves the capabilities of an LSTM network, its architecture remains more streamlined (Gu et al., (2019); Mahmoud et al., (2021); Ma et al., (2022)).
Specifically, this component accommodates the feature sequence over past time windows $1:t$, thereby learning temporal dependencies. We represent the temporal embedding as $\mathcal{Z}_{T}$, taking into account both historical spatiotemporal road features and traffic accidents as input:
@@ Appendix B Temporal encoder – GRU
Appendix C Spatial encoder – GAT

### v4

@@ 4.6.1 Temporal Stability of Metrics
Evaluating temporal stability is crucial for evaluating a model’s performance over multiple prediction time steps, particularly in data-driven methods, where superior results in specific instances might mask potential underfitting issues (Ding et al.,, 2022; Mannering,, 2018). Our primary goal is to achieve consistently low uncertainty, as indicated by the MPIW metric in three boroughs, depicted in Figure 4. We strive to maintain a narrow yet comprehensive MPIW throughout the prediction steps to ensure reliable prediction intervals for traffic crashes.
@@
As indicated in Figure 4, the MPIW is typically lowest on the first day of testing and generally increases over time. However, this increase is smoother in our STZITD-GNNs model compared to the more erratic fluctuations observed in baseline models. This smoother increase reflects our model’s ability to handle the inherent uncertainties and the variability of risk levels, which might be less accurately captured by models with misspecified distributions. Moreover, boroughs like Westminster and Lambeth, known for higher instances of zero-crash risk values, show more fluctuating MPIWs, suggesting greater uncertainties in prediction intervals due to the sparsity of traffic crashes. Despite these challenges, our model consistently delivers stable and accurate predictions across different boroughs.
In the evaluation of the temporal stability for the estimation of points, depicted in Figure 5, the STZITD-GNN customised approach emphasises the advantages of applying the most appropriate probabilistic assumptions, enhancing spatiotemporal representations, and managing the various aspects of traffic crashes. It is apparent that the STG-GNN model consistently shows the weakest performance across all temporal prediction periods, with deterministic models also performing poorly. Furthermore, models based on the Negative Binomial (NB) distribution exhibit unsatisfactory performance characterised by higher fluctuation and higher error rates. This inadequacy stems from the focus of NB models on binary crash-occurrence outcomes, which often overlook the complexities of long-tail crash risk values. In contrast, our model exhibits significantly less variance across all three evaluated metrics at each time step, demonstrating stability even toward the end of the prediction horizon. This robust performance is attributed to the model’s ability to dynamically fine-tune prediction intervals using four critical parameters, effectively addressing the challenges posed by the sparsity and skewed distribution of traffic crash data.
@@
In terms of the zero rate, deterministic deep learning models tend to conflate no-risk with lower-risk values. This occurs because deterministic models do not adequately differentiate between the absence of risk and the presence of minimal risk. As a result, these models often misclassify low-risk scenarios as zero-risk, particularly in situations with extreme zero inflation, leading to overfitting. For the hit rate, the issue is inversely presented as low-risk predictions often being incorrectly classified as no-risk. This misclassification is due to the model’s inability to capture subtle variations in risk levels.
The final component of our evaluation focusses on quantifying both non-crash roads and high-risk roads where actual crashes occurred, as shown in Figure 6. The STZITD-GNN model excels consistently in both metrics, maintaining stable performance with less than 18% variance throughout the 14-step period. Unlike deterministic models, which exhibit less variability but struggle significantly with accurately identifying true noncrash roads, often misclassifying all roads as having no crash risk, the STZITD-GNN model effectively addresses these challenges.
In scenarios of extreme zero inflation, deterministic models frequently misclassify roads, conflating no-risk with lower risk values, thereby leading to overfitting. Conversely, probabilistic models using Gaussian distributions struggle to manage the high prevalence of zero values and lower risk values, resulting in poor performance and high sensitivity in the ZR metrics over time. This problem occurs because each time step may present unique scenarios of zero-valued roads, which Gaussian models often misinterpret due to their uniform mean and variance assumptions.
In the identification of actual crashes on high-risk crash roads, as illustrated in Figure 6(b), the ZITD distribution performs exceptionally well. It employs a weighted dispersion parameter ($\phi$) that effectively distinguishes lower crash risks from non-crash roads, accurately reflecting the distribution of low, medium, and high-risk roads where actual crashes occurred, shown as long-tailed data distributions. This adaptability to the complex temporal dynamics of road-level crash data at each time step not only mitigates overfitting for non-crash scenarios, but also stabilises the model performance, enabling it to reliably capture real crash occurrences in critical geographic locations.
@@
(b) Accuate Hit Rate on Crash Occurrence at Top 20% Risky Roads over Time Performance
Figure 6: Time-wise Comparison for the Non-crash Roads Quantification and Accurate Crash Hit Metrics
4.6.2 Analysis of Long-tail Crash Data Characteristics
To explore the impact of inherently long-tailed data on our model, we created a three-dimensional surface plot over four information sources, as illustrated in Figure 7. Due to visualisation constraints, it is not feasible to represent all four parameters in a single plot. We therefore focus on the TD parameters $\phi$, $\rho$, and $\mu$, with the understanding that the sparsity parameter $\pi$ significantly influences the original TD parameters. This plot illustrates the predicted crash values, with the actual crash values $y$ represented by the colour gradient.
A distinct trend is evident in the learnt values of $\rho$, consistently approaching 2, suggesting a preference for more skewed distributions as $\rho$ increases from 1 to 2 (Tweedie et al.,, 1984; Kurz,, 2017). This trend corresponds with the long-tailed nature of our data and validates the probabilistic assumptions. The increase in $\rho$ serves as an adaptive response to the component of the loss function $y\frac{\mu^{1-\rho}}{1-\rho}(y>0)$, effectively penalising the underestimation of true, although infrequent, higher values of crash risk within the long tail. This adjustment directs the model to focus more on such kinds of traffic crash rates, continually benefiting from significant enhancements in model sensitivity and accuracy.
Additionally, significant variations in the values $\phi$ are observed, particularly concerning data with zero value, which serve as the dispersion parameter. For example, in Westminster, $\phi$ is significantly lower than in other boroughs, indicating fewer zero-inflation issues. Traditionally, $\phi$ in the TD distribution is crucial for managing the frequency of zero occurrences to effectively address zero inflation. However, due to the highly imbalanced nature of our data, a single $\phi$ parameter often fails to capture the prevalence of zeros, as demonstrated by the overlap of plots in areas of lower and zero crash values. In our ZITD model, an enhanced $\phi$, augmented by the sparsity parameter $\pi$, manages these features, offering a precise depiction of the inherent sparsity and discrete uncertainty in road-level crash predictions.
Moreover, the model parameter $\mu$, which denotes the mean of the predicted crash values, shows a strong correlation with the actual crash values $y$, as evidenced by the colour alignment in the graphs. This correlation not only substantiates the effectiveness of the model in estimating mean points, but also affirms its capability to accurately mirror and adapt to the complexities of urban crash data.
@@ 4.6.2 Analysis of Long-tail Crash Data Characteristics
Figure 7: Three Tweedie parameter visualization on Lambeth, Tower Hamlet, and Westminster with multi-step predict. The colour bar shows the predicted crash from the highest to zero.
@@ 5 Conclusion and Discussion
This paper introduces STZITD-GNN, a probabilistic deep learning framework that, to the best of our knowledge, is the first uncertainty-aware model specifically tailored for road-level crash prediction. Developed from data-driven insights, the STZITD-GNN utilises predefined distribution characteristics of traffic crash data to enhance the representation of sparse traffic crash scenarios. It effectively detects finer-level correlations in traffic crashes, thereby delivering precise predictions with narrow prediction intervals and reduced uncertainties. Through systematic multi-timescale experiments conducted on three real-life urban crash datasets, our results demonstrate that the efficacy of our STZITD-GNN approach surpasses that of state-of-the-art probabilistic deep learning models as well as deterministic models. The key contributions of this study are summarised as follows.
@@
  • Innovative Distribution Extension: We have adapted the conventional TD to a ZITD distribution, introducing a sparsity parameter, $\pi$, integrated with the dispersion parameter $\phi$. This modification tailors the model to our data-driven assumptions, representing the dispersed and extreme zero-inflated nature of crash data at the road level.
@@
  • Integration with ST-GNN: The incorporation of the ZITD encoder into the ST-GNN framework makes the STZITD-GNN model an end-to-end framework to handle the complex spatiotemporal dynamics of crashes. It surpasses traditional point estimation methods by enabling the mapping of output variance and prediction intervals. Additionally, the model facilitates an interpretable, uncertainty-aware process that elucidates the underlying factors contributing to improved prediction accuracy.
@@
  • Empirical Validation Across Boroughs: The model has shown improved predictive accuracy and minimal uncertainties compared with benchmark models in three London boroughs, each featuring unique spatial crash patterns in multi-step scenarios. This performance confirms the robustness and generalizability of the model across diverse urban settings, even with sparse datasets. Additionally, the model’s stable performance across heterogeneous datasets, which vary in crash determination criteria, further underscores its reliability and effectiveness in urban traffic risk assessment.
@@
It should be recognised that the challenges of overdispersion and zero inflation extend beyond traffic crashes, affecting a wide array of urban issues such as crime rates. This universality suggests that our proposed model could be applicable to a broader range of contexts beyond its current usage.
@@
While our research demonstrates its effectiveness in predicting traffic crashes, there are several limitations that provide direction for future research. We aim to enhance the accuracy of mapping crashes at finer temporal resolutions and plan to expand the model’s applicability to various countries under consistent assumptions. Future developments will focus on creating a more adaptable model that can accommodate geographic differences, thus increasing its usability across diverse regions. In addition, we intend to integrate more diverse data types into our model. Incorporating GPS data, for example, could provide more detailed insight into traffic exposure. To further improve the accuracy and utility of our model, we also plan to use more comprehensive data sets, including urban characteristics derived from street or satellite images. This would enrich our model’s contextual understanding and could identify factors that exacerbate traffic crash impacts across different locations and times. Addressing these challenges in future studies will improve the robustness and applicability of the framework in real-world scenarios, ultimately contributing to the development of more effective traffic management systems and crash prevention strategies.
@@
In conclusion, the development of the STZITD-GNN model marks a significant advancement in the prediction of urban road-level traffic crashes. However, this is just the beginning of an ongoing effort to refine and enhance this model. By integrating additional data sources and continually improving our methodologies, our goal is to develop a tool that not only mitigates traffic risks but also significantly contributes to global urban safety strategies.
@@ Appendix A Tweedie Table
Appendix B Tweedie Parameters
The frequency of traffic crashes $C_{k}$ adheres to a Poisson distribution represented by $Pois(\lambda)$ with a mean of $\lambda$, expressed as $C_{k}\sim Pois(\lambda)$. This is natural, as frequency often appears as the main prediction target in the previous literature (Zhou et al., 2020b, ; Trirat et al.,, 2023; Bonera et al.,, 2024). Currently, the severity of the crash $l_{k}^{(j)}$ is modelled as independent and identically distributed $i.i.d$ Gamma random variables, $C_{k}\perp l_{k}^{(j)}$ and $\forall j$, denoted as $Gamma(\alpha,\gamma)$. These variables have a mean of $\alpha\gamma$ and a variance of $\alpha\gamma^{2}$, such as $l_{k}^{(j)}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Gamma}(\alpha,\gamma)$. $C_{k}$ and $l_{k}^{(j)}$ together form the model of the occurrence of a single crash. Consequently, the traffic crash risk $y_{k}$ is formulated as a Poisson sum of these i.i.d. Gamma random variables. Considering Equation 1 and considering $21<\rho<$ 2, we can finally reformulate the random variable $y_{k}$ as follows:
In the scenario where $C_{k}=0$, we correspondingly have $y_{k}=0$. The probability density at zero for traffic crash risk, therefore, becomes $\mathbb{P}(y_{k}=0)=\mathbb{P}(C_{k}=0)=\exp(-\lambda)$. In contrast, when $C_{k}>0$, $y_{k}$ is deduced from the sum of $C_{k}$ independent Gamma random variables. For clarity, following the work of Zhou et al., 2022a , we re-parameterize the TD distribution, setting $\theta=\mu^{1-\rho}/(1-\rho)$ and $\kappa(\theta)=\mu^{2-\rho}/(2-\rho)$. This results in the revised TD distribution as:
where
In this case, $\mu,\phi$ and $\rho$ are the three parameters that govern the probability and expected value of the risk of traffic crashes. Parameters in the Gamma and Poisson distributions ($\lambda,\alpha,\gamma$) can be obtained as (Kurz,, 2017; Halder et al.,, 2019; Mallick et al.,, 2022), :
Appendix C Temporal encoder – GRU
Traffic crashes demonstrate a significant temporal correlation, underscoring the idea that the dynamic temporal features of the roads encapsulate a wealth of information. This observation has been acknowledged and exploited in prior methodologies (Wang et al., 2021a, ; Zhao et al.,, 2019). As shown in Figure 1, to seize these temporal proximity features, we implement a GRU (Chung et al.,, 2014). The GRU, in comparison to traditional LSTM networks, offers the advantage of learning relatively long-term dependencies without facing issues of vanishing and exploding gradients. In the GRU framework, there isn’t a segregation of internal and external states, as seen in LSTM networks. Instead, it addresses the gradient vanishing and exploding problem by directly integrating a linear dependency between the current network state $h_{t}$ and the preceding state $h_{t-1}$. Consequently, while a GRU preserves the capabilities of an LSTM network, its architecture remains more streamlined (Gu et al., (2019); Mahmoud et al., (2021); Ma et al., (2022)).
Specifically, this component accommodates the feature sequence over past time windows $1:t$, thereby learning temporal dependencies. We represent the temporal embedding as $\mathcal{Z}_{T}$, taking into account both historical spatiotemporal road features and traffic crashes as input:
@@ Appendix C Temporal encoder – GRU
Appendix D Spatial encoder – GAT